A cell that consists of a standard hydrogen half-cell and a Cu/Cu2+ half-cell has Eo = 0.34 V. If the cell contains 0.00318 M concentrations of HCl and CuSO4, what potential does it produce?
Answer #1

Cu2+ + 200 ils) E= 0.34 E = E + 0.059 log [427] ชา = 0.34 + 0.059 log (0.00318) - 0.34 + 0,059* (-2.5) 2 - 0.34 - 0.07 = 0.26